OpenStudy (anonymous):

1. A group holds a raffle to raise money. They sell 100 raffle tickets for $5 each. There is one grand prize worth $100, three second place prizes worth $25 each and five 3rd place prizes worth $10 each. Find the expected value associated with the purchase of one ticket. (Remember to subtract the cost of playing from your winnings)

OpenStudy (anonymous):

anyone can help me?

OpenStudy (anonymous):

HELP ME PLZZZ

OpenStudy (valpey):

Okay, how much do they collect in ticket sales, and how much in total do they pay out in prizes?

OpenStudy (anonymous):

500 AND 225?

OpenStudy (valpey):

Yes, they collect $500 in ticket sales but they must pay out prizes of $100 plus 3*$25 = $75 plus 5*$10=$50 or $225 in total prizes.

OpenStudy (valpey):

So the gross expected value of a ticket is the total prize pool divided by the number of tickets. The net expected value (what the question is asking) is that amount minus the cost of buying a ticket.

OpenStudy (anonymous):

so the answer is 2.25?

OpenStudy (valpey):

The people buying the tickets aren't expecting to profit (unless they have unrealistic expectations about luck or something). They are choosing a fun way of donating money to the Group. They are expecting to lose money, or we can say the expected value of a ticket is negative. The raffle is a fund raiser and the person purchasing the ticket can think of $2.75 going toward the Group and the $2.25 going toward the prize pool.

OpenStudy (valpey):

But in terms of the answer to the question, it is the amount that a person buying a ticket expects to lose, so $2.25 - $5.00.

OpenStudy (anonymous):

Oh... I still confused about it... Do I need to use the probability to calculate it? Such as I have \[\frac{ 1 }{ 100 }\] chance to get $100?

OpenStudy (valpey):

You could do it that way for each prize, but it will end up the same.

OpenStudy (valpey):

You will see that \(\frac{1}{100}*$100+\frac{3}{100}*$25+\frac{5}{100}*$10=$2.25\)

OpenStudy (anonymous):

I gotcha! Thank you so much!